Working together, Brad and Mike can complete a job in 3 hours. If it takes mike 5 hours to do the job alone,how long would it take Brad to do the job alone?
PLZ help

Answer by stanbon(75887)

Working together, Brad and Mike can complete a job in 3 hours. If it takes mike 5 hours to do the job alone,how long would it take Brad to do the job alone?
------
Together rate:: 1/3 job/hr
----------
Mike rate: 1/5 job/hr
----
Brad rate: 1/x job/hr
==========================
Equation:
rate + rate = together rate
1/x + 1/5 = 1/3
----
15 + 3x = 5x
2x = 15
x = 7.5 hrs (time for Brad to do the job alone)
===============================================
Cheers,
Stan H.
=============